Q1. Catalase is an important enzyme that breaks down its substrate, H2O2, into water and oxygen in nearly all living cells. The enzyme has k2 = 1.1*107 sec-1 and Km = 2.0 mM. An experiment was set up under the conditions when the reaction rate is independent of [H2O2]. In this experiment, at what concentration of catalase (use SI units and scientific notation) will the reaction have a rate of 850 mM/min?


Q2. The degradation of ethyl acetate (EA) is a first order process with t1/2 = 110 min.

a) Fill in the blank: For this reaction, a graph of _______________ vs. time will give a straight line (linear plot).

b) What is the rate constant for the degradation of EA ?

c) What percentage of EA will remain after 440 minutes?


1
Expert's answer
2022-02-02T04:05:03-0500

Question 1)

[A]/[n]n=kt.

Where [A] is the concentration of the reactants at time t.

[n] n is the concentration of the reactants at t=0.

K is the rate constant.

Therefore:

1.1×107sec-¹×2.0=850/[n]n.

=850/1.1×107×0.2





Question 2)

a)the graph of degradation of ethyl acetate vs time will give a straight line.

b)t½=110min

K=0.693/110min=6.3×10-³

=0.693/110×60=0.693/6600sec=1.05×10-⁴sec-¹

c)After 4 half lives, percentage of EA will remain

=100%/2⁴

=100%/16

=6.25%